OpenStudy (anonymous):

Given f(x)= square root of x+8 and g(x)=-4/x find the following: (f*g)(1) the * is actually a little circle I couldn't find a way to type it in

OpenStudy (anonymous):

g(1)= -4/1 = -4 f(g(1))= sqrt -4 + 8 = sqrt 4 = 2